# arXiv:2603.00637v1[cond-mat.mtrl-sci] 28 Feb 2026

### v1

@@ IV.2 Semiconductors and Insulators
We further observe that the bands immediately below the valence-band maximum are very similar between QSGWGW and $G^{0}W^{0}$@PBE, whereas deeper valence bands in the energy region around -17 eV are shifted to lower energies in QSGWGW (Fig. 6). This non-rigid re-alignment reflects the state- and energy-dependent nature of the $GW$ self-energy and the accompanying changes in screening along the self-consistent cycle, as also discussed in the literature Kotani et al. (2007); Salas-Illanes et al. (2022). We expect that the energy positions of these deeper valence states are better described in QSGWGW.
In summary, benchmark calculations demonstrate that our all-electron QSGWGW implementation in LibRPA is reliable and accurate. It exhibits the typical systematic overestimation of band gaps characteristic of the QSGWGW approximation, consistent with other state-of-the-art implementations. The combination of NAOs, LRI, and the space-time formalism provides a practical all-electron framework for QSGWGW calculations for both molecules and periodic solids within LibRPA.
@@ V Conclusion and outlook
We have developed and validated an all-electron implementation of the QSGWGW method for molecules and periodic systems within the numerical atomic orbital (NAO) framework of LibRPA, based on the space-time formalism and the localized resolution of identity (LRI) technique.
We analyzed numerical aspects specific to QSGWGW in this framework, in particular, the stability of the Padé analytic continuation and its basis representation dependence, and identified “Mode B” as the more robust construction within our current protocol. Benchmark calculations for a set of small molecules and for representative semiconductors and insulators show that the resulting IPs and quasiparticle gaps are consistent with established reference implementations, providing strong validation of the correctness of our implementation. For the crystalline solids, we obtain an MARE of 9.43% (MPE of +9.43+9.43%) relative to experiment, while for the molecular subset in Table 1, we obtain an MARE of 3.64% (MPE of +2.12+2.12%) for the vertical ionization potentials relative to experiment.
Consistent with the known behavior of standard QSGWGW, the gaps are systematically overestimated, reflecting under-screening of the Coulomb interaction in the absence of vertex corrections van Schilfgaarde et al. (2006); Bruneval and Gatti (2014).

### v2

@@ IV.2 Semiconductors and Insulators
We further observe that the bands immediately below the valence-band maximum are very similar between QSGWGW and $G^{0}W^{0}$@PBE, whereas deeper valence bands in the energy region around -17 eV are shifted to lower energies in QSGWGW (Fig. 6). This shift reflects the state- and energy-dependent nature of the $GW$ self-energy and the accompanying changes in screening along the self-consistent cycle, as also discussed in the literature Kotani et al. (2007); Salas-Illanes et al. (2022). We expect that the energy positions of these deeper valence states are better described in QSGWGW.
To wrap up this section, benchmark calculations demonstrate that our all-electron QSGWGW implementation in LibRPA is reliable and accurate. It exhibits the typical systematic overestimation of band gaps characteristic of the QSGWGW approximation, consistent with other state-of-the-art implementations. The combination of NAOs, LRI, and the space-time formalism provides a practical all-electron framework for QSGWGW calculations for both molecules and periodic solids within LibRPA.
@@ V Conclusion and outlook
In summary, We have developed and validated an all-electron implementation of the QSGWGW method for molecules and periodic systems within the numerical atomic orbital (NAO) framework of LibRPA, based on the space-time formalism and the LRI technique.
We analyzed numerical aspects specific to QSGWGW in this framework, in particular, the stability of the Padé analytic continuation and its basis representation dependence, and identified “Mode B” as the more robust construction within our current protocol. Benchmark calculations for a set of small molecules and for representative semiconductors and insulators show that the resulting IPs and quasiparticle gaps are consistent with established reference implementations, providing strong validation of the correctness of our implementation. For the crystalline solids, we obtain an MARE of 9.43% (MRE of +9.43+9.43%) relative to experiment, while for the molecular subset in Table 1, we obtain an MARE of 3.64% (MRE of +2.12+2.12%) for the vertical ionization potentials relative to experiment. Consistent with the known behavior of standard QSGWGW, the gaps are systematically overestimated, reflecting under-screening of the Coulomb interaction in the absence of vertex corrections van Schilfgaarde et al. (2006); Bruneval and Gatti (2014).
@@
Last but not least, our present QSGWGW is built upon the existing low-scaling $G^{0}W^{0}$ implementation in LibRPA, which enables efficient $G^{0}W^{0}$ calculations for systems containing up to $O(10^{3})$ atoms Zhang et al. (2026). This low-scaling algorithm directly carries over to QSGWGW, and benchmark calculations for its performance on large-scale systems are ongoing. We expect that our QSGWGW implementation will provide a highly useful tool for investigating complex materials systems.
Data availability
The data that support the findings of this study are available upon reasonable request from the authors.